OpenStudy (anonymous):

Multiple Choice: An equation for a tangent to the graph of y=archsin (x/2) at the origin is (a) x-2y=0 (b) x-y=0 (c) x=0 (d) y=0 (e) pix-2y=0

OpenStudy (anonymous):

just arcsin* btw.. hehe

OpenStudy (anonymous):

find the derivative then substitute 0,0

OpenStudy (anonymous):

derivative of arcsin is 1/(sqrt 1-x^2) btw..

OpenStudy (amistre64):

y = sinh^-1 y' = e^-x - e^...../2........ something..... I know the other ones :)

OpenStudy (amistre64):

oh...that "h" threw me. y' = 1/sqrt(1-x^2)

OpenStudy (anonymous):

lol !

OpenStudy (anonymous):

thats why i corrected him at top :P

OpenStudy (amistre64):

the equation to the tangent has the derivative as the slope.... 0 = f'(x)x + b

OpenStudy (amistre64):

yeah ;)

OpenStudy (amistre64):

"y=archsin (x/2) at the origin" ack!!..... it helps to read the whole thing, I get lazy and start making up my own question lol

OpenStudy (amistre64):

at 0,0 you get a slope of 1, so the equation of the tangent is: y = x and thats my final offer

OpenStudy (anonymous):

lmao! i hope u put 0,0 in the derivative!

OpenStudy (amistre64):

:) I think I miss placed a Dx tho.... since its got that x/2 in there ;)

OpenStudy (anonymous):

\[1/ (\sqrt{2-x} \sqrt{x+2}\]

OpenStudy (amistre64):

\[y' = \frac{Dx}{\sqrt{1-\frac{x}{2}^2}} \rightarrow Dx = \frac{1}{2}\]

OpenStudy (anonymous):

thats the derivative....forgot to close bracket lol

OpenStudy (anonymous):

I DID IT DIFF! :P

OpenStudy (amistre64):

I wanna change my vote :) y = x/2 ;) but I probably missed that one too

OpenStudy (anonymous):

lol! ha this is bad... i dont like this question!

OpenStudy (amistre64):

ack!! how can a fraction = 0 if its got a constant up top....

OpenStudy (anonymous):

yo this is messed...

OpenStudy (anonymous):

cuz the derivative of that equation will be the equation of the tangent slope...it cant be ny of those options :S...

OpenStudy (amistre64):

take this question, throw it in a sack, burn it in the fireplace, kill a chicken and spread its blood around the place to ward off any evil spirits that might rise from the ashes... ;)

OpenStudy (amistre64):

come to think of it, when does sin^-1 have a hump in it?

OpenStudy (anonymous):

lmao!!!

OpenStudy (anonymous):

I thank you guys for all of the help! I can't exactly figure it out myself. Tricky tricky. tisk tisk. :)